Please help me answer the following homework problem. I do not understand how to solve it.
Given f(x)= 1-5x and g(x)= x^2-3 find the composite function and simplify.

Answer by stanbon(75887)

Given f(x)= 1-5x and g(x)= x^2-3 find the composite function and simplify.
--------------------------------
f[g(x)] = f[x^2-3] = 1 - 5(x^2-3) = 16 -5x^2
-------------------
g[f(x)] = g[1-5x] = (1-5x)^2-3 = 1-10x+25x^2-3 = 25x^2-10x-2
==========================
Cheers,
Stan H.